OpenStudy (anonymous):

Solve cos(x)(cosx+1)=0 @terenzreignz :)

OpenStudy (anonymous):

answer choices A,B,C,D from top to bottom :) **I'm not too sure, but I'm leaning towards A... what do you think? :)

terenzreignz (terenzreignz):

Same technique..., even though it's not really necessary, replace cos(x) with u u(u+1) = 0 Possible values for u?

OpenStudy (anonymous):

0, -1 ?

terenzreignz (terenzreignz):

Yup.... so cos(x) =0 cos(x) =-1 Possible values for x?

OpenStudy (anonymous):

pi, pi/2, 3pi/2, and 2pi ?

OpenStudy (anonymous):

or 90, 180, 270, 360 degrees?

terenzreignz (terenzreignz):

2pi? \[\Large \cos(2\pi) = 1\]

OpenStudy (anonymous):

yup! 1=1 :)

terenzreignz (terenzreignz):

We don't have a \(\large \cos(x) = 1\)

OpenStudy (anonymous):

im not following :( what do u mean?

terenzreignz (terenzreignz):

\(2\pi\) is not a solution... Because we only have cos(x) = 0 and cos(x) = -1 We do NOT HAVE cos(s) = 1

OpenStudy (anonymous):

okay... so whats next then? :/

terenzreignz (terenzreignz):

cos(x) *

terenzreignz (terenzreignz):

What's next? You have not even finalised your answer for the current step... What are the values for x which would make cos(x) = 0 or cos(x) = -1 true?

OpenStudy (anonymous):

lol yeah haha i didn't mean like the next step haha i meant like how can i figure this part out.. :P lol sorry bout that! :P so do i plug pi, pi/2, 3pi/2 into cos(x) and see if they equal 0 or -1? and not 2pi tho right?

terenzreignz (terenzreignz):

yeah, so, are the answers \(\Large \pi \ , \ \frac{\pi}{2} \ , \frac{3\pi}{2}\) correct?

OpenStudy (anonymous):

i believe so... they are, aren't they? :)

terenzreignz (terenzreignz):

I'm asking you. I'm asking you to be sure.

OpenStudy (anonymous):

cos(pi)= -1 cos(pi/2) = 0 cos(3pi/2) = 0 so yeah right? sorry i probs should have included this part in the thing above lol :P

terenzreignz (terenzreignz):

Okay. Good enough. So your final answer?

OpenStudy (anonymous):

B ?

terenzreignz (terenzreignz):

Yeah. Current Streak : 0 Best Streak : 3 You were leaning towards A... so... nope :P

OpenStudy (anonymous):

hahah yeah :( back to 0 :( lol but yay! :) this one makes sense now tho!!! :) thanks!! :)